OpenStudy (anonymous):

what is the common ratio for the geometric sequence 1/128, 1/16, 1/2, 4

OpenStudy (jango_in_dtown):

divide the 1st term by 2nd term

OpenStudy (jango_in_dtown):

\[\frac{ 1 }{ 128 }\div \frac{ 1 }{ 16 }=\frac{ 16 }{ 128 }=1/8\]

OpenStudy (anonymous):

so it would be 1/8?

OpenStudy (jango_in_dtown):

yeah

OpenStudy (anonymous):

i thought it would had been -4

OpenStudy (jango_in_dtown):

no.. for germetric series common ratio is given by \[\frac{ a_2 }{ a_1 }=\frac{ a_3 }{ a_2 }=.....\]

OpenStudy (anonymous):

ooohhhh ok thanks alot :D

OpenStudy (jango_in_dtown):

now you check that the result is true by dividing the third term by 2nd term